# GenoTEX: A Benchmark for Automated Gene Expression Data Analysis in Alignment with Bioinformaticians

### v2

@@ Abstract
Recent advancements in machine learning have significantly improved the identification of disease-associated genes from gene expression datasets. However, these processes often require extensive expertise and manual effort, limiting their scalability. Large Language Model (LLM)-based agents have shown promise in automating these tasks due to their increasing problem-solving abilities. To support the evaluation and development of such methods, we introduce GenoTEX, a benchmark dataset for the automated analysis of gene expression data. GenoTEX provides annotated code and results for solving a wide range of gene identification problems, encompassing dataset selection, preprocessing, and statistical analysis, in a pipeline that follows computational genomics standards. The benchmark includes expert-curated annotations from bioinformaticians to ensure accuracy and reliability. To provide baselines for these tasks, we present GenoAgent, a team of LLM-based agents that adopt a multi-step programming workflow with flexible self-correction, to collaboratively analyze gene expression datasets. Our experiments demonstrate the potential of LLM-based methods in analyzing genomic data, while error analysis highlights the challenges and areas for future improvement. We propose GenoTEX as a promising resource for benchmarking and enhancing automated methods for gene expression data analysis. The benchmark is available at https://github.com/Liu-Hy/GenoTex.
@@ 1 Introduction
Despite the scientific value of gene data analysis, these tasks are often repetitive, labor-intensive, and prone to errors [8]. The rapid increase in transcriptomic data and potentially inefficient workflows lead to considerable financial burden [40]. The genetics research industry incurs an annual expense of around \$848.3 million on manual data analysis tasks [63], with costs expected to increase at a compound annual growth rate (CAGR) of 12% [63] to 16% [64] by 2030. Bioinformaticians spend significant effort on these repetitive tasks, valued at around \$29 per hour [60]. This high volume of routine tasks greatly impacts job satisfaction among bioinformatics professionals, as surveys show that data scientists, including bioinformaticians, prefer engaging in advanced analytical tasks rather than routine data processing [84]. Currently, up to 45% of their work hours are spent on tasks that could be automated [84]. These financial and workforce challenges highlight the urgent need for more efficient and cost-effective data analysis solutions in genetics research [6].
@@
To facilitate the evaluation and development of such methods, we propose the Genomics Data Automatic Exploration Benchmark (GenoTEX), a benchmark dataset for the automated analysis of gene expression datasets to identify disease-associated genes while considering the influence of other biological factors. Following the standards of computational genomics, and based on the common practices of skilled human bioinformaticians, we unified the process of analyzing various gene expression datasets for solving gene identification problems into a standardized pipeline, with detailed procedures documented in a guidelines file (Appendix B). We then trained and organized a group of bioinformaticians to manually perform the data analysis according to these guidelines, creating a benchmark dataset consisting of input data, annotated code, and intermediate and final analysis results. Fig. 1 illustrates our benchmark curation pipeline. Based on this benchmark, we propose three tasks, namely, dataset selection, data preprocessing, and statistical analysis, with corresponding metrics for evaluating different aspects of the analysis process.
@@
Furthermore, to provide baselines for these tasks, we propose GenoAgent, a team of LLM-based agents that simulate the behavior of bioinformaticians in gene data analysis. GenoAgent adopts a multi-agent collaboration pattern that facilitates task decomposition and code correction, to perform the end-to-end data analysis for solving gene identification problems. Each programming agent is instructed with detailed guidelines to perform tasks in a Jupyter Notebook-style multi-step workflow.
@@
  • We propose a benchmark dataset, GenoTEX, that presents the analysis pipeline for a rich set of gene identification problems, with documented code and output. We believe it will serve as a useful resource for the evaluation and development of advanced methods for automated gene expression data analysis.
@@ 2 Related work
Our work takes a different approach by leveraging state-of-the-art LLMs without additional domain-specific training. Instead, we develop structured prompting techniques and communication protocols that enable LLM-based agents to perform sophisticated planning, analysis, and coding tasks required for genomic data exploration. This approach demonstrates how general-purpose LLMs can be effectively applied to specialized scientific domains through careful system design rather than model customization.
@@ 3.1 Standardized pipeline of gene expression data analysis
Evaluating the automatic exploration of this kind of problems is complex due to its nature. The combination of different traits and conditions leads to a multitude of gene identification scenarios, many of which remain understudied in biomedical literature. Moreover, while data-driven approaches provide valuable insights, they must ultimately be combined with interventional biological experiments or clinical trials to confirm the significance of identified genes. This absence of a clear “ground truth” complicates the evaluation of our analysis results. Defining the exact insights that should be extracted from our data analysis is therefore complex.
@@ 3.1.1 Data preprocessing
This step aims to filter out irrelevant datasets and select the best dataset (or pair of datasets) to analyze for each gene identification problem. The selection of datasets involves three steps:
@@
Figure 2: The pipeline for preprocessing a GEO series dataset.
@@ 3.1.2 Statistical analysis
To ensure reliable gene identification, the pipeline incorporates steps to correct potential confounding factors [11], including batch effects [44]. One important type of confounding arises when the distribution of gene expressions varies across subgroups within the data due to sample population heterogeneity rather than the disease itself [92]. This variation introduces significant bias, as spurious gene-disease associations may be detected due to technical or population differences rather than genuine biological relationships [76]. We directly use a Python software package developed to handle different scenarios of confounding factors [76].
@@ 3.2 Benchmark creation
Figure 3: Descriptive statistics of our GenoTex benchmark.
Gene Identification Problems
@@
Lines of code for analyzing per dataset 92$\pm$28
Total lines of code for analysis 84,023
Normalized gene features per dataset 15246 $\pm$ 4233
Significant genes identified per problem 47$\pm$62
Gene identification problem design
We began by curating a scientifically relevant list of human traits spanning from major public health concerns to representative rare diseases, along with physical characteristics like height and bone density. A computational biologist compiled 132 traits across 9 main categories, including cardiovascular diseases and neurological disorders. This yielded our first set of problems: What are the significant genes related to the trait? (unconditional gene identification).
We then created trait-condition pairs by matching each trait with a condition—either another trait from our list or demographic attributes like age and gender. From the 17,556 possible trait-condition pairs, we selected the most scientifically valuable ones to frame our second problem set: What are the significant genes related to the trait when considering the influence of the condition? (conditional gene identification).
Our selection of trait-condition pairs involved both domain expertise and data-driven approaches. We applied manually designed rules determining which pairs to include or exclude based on trait categories (detailed in Appendix D). For undecided pairs, we measured trait-condition association by calculating Jaccard similarity between gene sets related to each trait and condition, using data from the Open Targets Platform [70]. We selected pairs with Jaccard similarity exceeding 0.1, as these likely share underlying genetic mechanisms, offering valuable insights into complex trait-condition interactions. This process yielded 1,252 conditional gene identification problems of significant scientific interest. Combined with our 132 unconditional problems, these constitute our benchmark’s complete problem set.
@@
To address our research problems, we sourced cohort datasets containing gene expression and clinical data from leading public repositories: (1) The Gene Expression Omnibus (GEO) [14], the largest available gene expression database; (2) The Cancer Genome Atlas (TCGA) [72], the leading cancer-focused gene expression database, accessed via the UCSC Xena platform [28]. Additionally, we acquired essential domain knowledge from the Open Targets Platform [70] for gene-trait associations and the NCBI Genes database [10] for gene synonym mapping. Appendix E provides detailed information about these data sources.
@@
Our team of 4 researchers designed the problem list and acquired relevant input data from public sources. During the pilot stage, a computational biologist and a doctoral student collaboratively developed example code for solving gene identification problems related to two traits. By extracting common patterns from these examples, they developed an initial version of the analysis guidelines for the entire benchmark. Then they iteratively refined the guidelines based on their experience analyzing 193 more problems, establishing a working framework for benchmark curation.
@@
To evaluate annotation consistency, we measured Inter-Annotator Agreement (IAA) between the two annotation versions. Results demonstrate high quality, with an F1 score of 94.73% for dataset filtering. For more evaluation of the benchmark quality, we take the analysis results in the final benchmark as the reference set, and computed the average performance of the two annotation versions using the metrics defined in Section 3.3. This also serves as a strong baseline of human expert performance, with results presented in Table 1. The set of selected genes is sensitive to specific choices made during cohort-specific feature encoding, where multiple reasonable approaches often exist. Despite this inherent ambiguity in gene selection, the AUROC score of 0.89 demonstrates high consistency among annotators, validating the reliability of our benchmark.
@@ 3.3 Tasks and metrics
Gene expression data preprocessing involves complex decisions that affect both attribute names and sample IDs. We evaluate preprocessing performance using three complementary metrics: (i) Attribute Jaccard (AJ) is the Jaccard similarity between sets of attributes of two datasets. It evaluates how well the method extracts attributes from the dataset by encoding clinical features and normalizing gene symbols. (ii) Sample Jaccard (SJ) is the Jaccard similarity between sets of sample IDs of two datasets. It measures how well the method integrates features from the same samples and handles missing values. Based on these metrics, we define (iii) Composite Similarity Correlation (CSC) as the product of the Attribute Jaccard, Sample Jaccard, and the Pearson correlation of common feature vectors (shared rows and columns) between datasets. This metric captures both structural and content similarity of datasets, so we consider it as the primary metric for evaluating preprocessing quality. We compute these scores per dataset and average across all datasets in our benchmark.
@@
Statistical analysis aims to identify trait-related significant genes. We evaluate this process using multiple complementary metrics. Treating gene identification as a binary classification task, we employ Precision, Recall, and F1 scores. The Area Under the Receiver Operating Characteristic curve (AUROC) [33] provides a threshold-independent measure of discrimination ability, particularly valuable for imbalanced gene sets. We also incorporate the Enrichment Score (ES) from Gene Set Enrichment Analysis (GSEA) [68], a standard metric in bioinformatics capturing the biological relevance of identified gene sets. We select AUROC as our primary metric due to its robustness to class imbalance and widespread acceptance in the machine learning community. We compute these scores per problem and average across all problems in our benchmark.
@@ 4.1 Motivation and agent design
We believe that integrating these components is essential for enabling agent systems to effectively tackle the complex task of gene expression data analysis. Thus, to propose reasonable baselines for our benchmark, inspired by human bioinformaticians’ workflow in gene data analysis, we propose GenoAgent, a team of LLM-based agents with specialized roles mirroring those in a genomic data science team. A Project Manager coordinates the analysis process for solving each gene identification problem, assigning tasks to agents with the standardized pipeline in our benchmark as instructions. Two programming agents, Data Engineer and Statistician, focus on the data preprocessing and statistical analysis tasks, respectively. When writing code to perform specific steps, they can access, adapt, and utilize function tools from a provided library file as needed. A Code Reviewer agent helps the programming agents debug code and verify that their code follows the instructions. A Domain Expert agent provides professional knowledge consultation to programming agents when required for data processing.
@@ 5 Experiment
This section describes our experiments to evaluate GenoAgent and other baseline methods on the GenoTEX benchmark. We conducted an end-to-end evaluation where methods process raw input data to complete the full analysis for solving gene identification problems. Additionally, we assessed these methods on each individual task to identify specific strengths and weaknesses. All evaluations use the tasks and metrics defined in Section 3.3. We tested GenoAgent on a wide range of representative LLMs, as shown in Table 1.
@@
Table 1: End-to-end performance of GenoAgent on various LLMs; ablation results of GenoAgent with OpenAI o1 model; human expert performance, and simple data-free baselines. We report performance on the gene identification problems in our benchmark, on the additional task of trait prediction, and runtime performance including success rate and efficiency metrics. Success rate (Succ.) refers to the percentage of problems for which the generated analysis code executes successfully. We use 00 as the trait prediction scores when code execution fails. “Tk.(K)” reports average input/output tokens (in thousands) per problem, while “Cost (\$)” and “Time (s)” represent average API cost and runtime per problem.
Model Gene Identification Trait Prediction Runtime Performance
@@
Model Gene Identification Trait Prediction Runtime Performance
@@ 5.1 Results
For comparison, we implemented a simple baseline where an OpenAI o1 model was directly prompted to identify significant genes for each problem. This approach yielded poor performance (0.55 AUROC), underscoring the difficulty of this task. Beyond gene identification, we also measured the agents’ ability to predict binary traits using gene expression features and other covariates from their selected datasets, with strong results confirming the validity of the regression models they implemented.
@@ 5.2 Discussions
While our results demonstrate the potential of LLM-based methods for gene expression analysis, they also highlight important limitations of current approaches.

### v3

# GenoTEX: An LLM Agent Benchmark for Automated Gene Expression Data Analysis

@@ Abstract
Recent advancements in machine learning have significantly improved the identification of disease-associated genes from gene expression datasets. However, these processes often require extensive expertise and manual effort, limiting their scalability. Large Language Model (LLM)-based agents have shown promise in automating these tasks due to their increasing problem-solving abilities. To support the evaluation and development of such methods, we introduce GenoTEX, a benchmark dataset for the automated analysis of gene expression data. GenoTEX provides analysis code and results for solving a wide range of gene-trait association problems, encompassing dataset selection, preprocessing, and statistical analysis, in a pipeline that follows computational genomics standards. The benchmark includes expert-curated annotations from bioinformaticians to ensure accuracy and reliability. To provide baselines for these tasks, we present GenoAgent, a team of LLM-based agents that adopt a multi-step programming workflow with flexible self-correction, to collaboratively analyze gene expression datasets. Our experiments demonstrate the potential of LLM-based methods in analyzing genomic data, while error analysis highlights the challenges and areas for future improvement. We propose GenoTEX as a promising resource for benchmarking and enhancing automated methods for gene expression data analysis. The benchmark is available at https://github.com/Liu-Hy/GenoTEX.
@@ 1 Introduction
Among the various approaches in computational biology, differential gene expression analysis identifies genes whose expression levels vary between biological conditions [49]. Our work focuses on gene-trait association (GTA) problems, which specifically examine how genes relate to traits (such as diseases) while accounting for additional biological factors or conditions [31, 53]. These GTA problems enable important clinical applications [79], where understanding gene expression patterns across diverse patient conditions can reveal context-dependent biomarkers and treatment targets [54].
Despite the scientific value of gene expression data analysis, these tasks are often repetitive, labor-intensive, and prone to errors [8]. The rapid increase in transcriptomic data and potentially inefficient workflows lead to considerable financial burden [41]. The genetics research industry incurs an annual expense of around \$848.3 million on manual data analysis tasks [67], with costs expected to increase at a compound annual growth rate of 12% [67] to 16% [68] by 2030. Bioinformaticians spend significant effort on these repetitive tasks, valued at around \$29 per hour [64]. This high volume of routine tasks greatly impacts job satisfaction among bioinformatics professionals, as surveys show that data scientists, including bioinformaticians, prefer engaging in advanced analytical tasks rather than routine data processing [89]. Currently, up to 45% of their work hours are spent on tasks that could be automated [89]. These financial and workforce challenges highlight the urgent need for more efficient and cost-effective data analysis solutions in genetics research [6].
@@
To facilitate the evaluation and development of such methods, we propose the Genomics Data Automatic Exploration Benchmark (GenoTEX), a benchmark dataset for the automated analysis of gene expression datasets to solve GTA problems while considering the influence of other biological factors. Following the standards of computational genomics, and inspired by the behaviors of skilled human bioinformaticians, we unified the process of analyzing various gene expression datasets for GTA analysis into a standardized pipeline, with detailed procedures documented in a guidelines file (Appendix B). We then trained and organized a group of bioinformaticians to manually perform the data analysis according to these guidelines, creating a benchmark dataset consisting of input data, annotated code, and intermediate and final analysis results. Fig. 1 illustrates our benchmark curation pipeline. Based on this benchmark, we propose three tasks, namely, dataset selection, data preprocessing, and statistical analysis, with corresponding metrics for evaluating different aspects of the analysis process.
@@
Furthermore, to provide baselines for these tasks, we propose GenoAgent, a team of LLM-based agents that simulate the behavior of bioinformaticians in gene data analysis. GenoAgent adopts a multi-agent collaboration pattern that facilitates task decomposition and code correction, to perform the end-to-end data analysis for solving GTA problems. Each programming agent is instructed with detailed guidelines to perform tasks in a Jupyter Notebook-style multi-step workflow.
@@
  • We propose a benchmark dataset, GenoTEX, that presents the analysis pipeline for a rich set of GTA problems, with documented code and output. We believe it will serve as a useful resource for the evaluation and development of advanced methods for automated gene expression data analysis.
@@ 2 Related work
Our work leverages state-of-the-art LLMs without additional domain-specific training. Instead, we develop structured prompting techniques and communication protocols that enable LLM-based agents to perform sophisticated planning, analysis, and coding tasks required for genomic data exploration. This approach demonstrates how general-purpose LLMs can be effectively applied to specialized scientific domains through careful system design rather than model customization.
@@ 3.1 Standardized pipeline of gene expression data analysis
Evaluating the automatic exploration of this kind of problems is complex due to its nature. The combination of different traits and conditions leads to a multitude of GTA problem scenarios, many of which remain understudied in biomedical literature. Moreover, while data-driven approaches provide valuable insights, they must ultimately be combined with interventional biological experiments or clinical trials to confirm the significance of identified genes. This absence of a clear “ground truth” complicates the evaluation of our analysis results. Defining the exact insights that should be extracted from our data analysis is therefore complex.
@@ 3.1.1 Data preprocessing
This step aims to filter out irrelevant datasets and select the best dataset (or pair of datasets) to analyze for each GTA problem. The selection of datasets involves three steps:
@@
Figure 2: High-level schematic of our GEO data preprocessing pipeline, with example code of core components that omits techinical details.
@@ 3.1.2 Statistical analysis
To ensure reliable GTA analysis, the pipeline incorporates steps to correct potential confounding factors [11], including batch effects [45]. One important type of confounding arises when the distribution of gene expressions varies across subgroups within the data due to sample population heterogeneity rather than the disease itself [98]. This variation introduces significant bias, as spurious gene-disease associations may be detected due to technical or population differences rather than genuine biological relationships [81]. We directly use a Python software package developed to handle different scenarios of confounding factors [81].
@@ 3.2 Benchmark creation
Table 1: Descriptive statistics of our GenoTex benchmark.
GTA Problems
@@
Lines of code for analyzing per dataset 261 $\pm$ 108
Total lines of code for analysis 237,907
Normalized gene features per dataset 18530 $\pm$ 6714
Significant genes identified per problem 359 $\pm$ 265
GTA problem design
We began by curating a scientifically relevant list of human traits spanning from major public health concerns to representative rare diseases, along with physical characteristics like height and bone density. A computational biologist compiled 132 traits across 9 main categories, including cardiovascular diseases and neurological disorders. This yielded our first set of problems: What are the significant genes related to the trait? (unconditional GTA problem).
We then created trait-condition pairs by matching each trait with a condition—either another trait from our list or demographic attributes like age and gender. From the 17,556 possible trait-condition pairs, we selected the most scientifically valuable ones to frame our second problem set: What are the significant genes related to the trait when considering the influence of the condition? (conditional GTA problem).
Our selection of trait-condition pairs involved both domain expertise and data-driven approaches. We applied manually designed rules determining which pairs to include or exclude based on trait categories (detailed in Appendix D). For undecided pairs, we measured trait-condition association by calculating Jaccard similarity between gene sets related to each trait and condition, using data from the Open Targets Platform [74]. We selected pairs with Jaccard similarity exceeding 0.1, as these likely share underlying genetic mechanisms, offering valuable insights into complex trait-condition interactions. This process yielded 1,252 conditional GTA problems of significant scientific interest. Combined with our 132 unconditional problems, these constitute our benchmark’s complete problem set.
@@
To address our research problems, we sourced cohort datasets containing gene expression and clinical data from leading public repositories: (1) The Gene Expression Omnibus (GEO) [14], the largest available gene expression database; (2) The Cancer Genome Atlas (TCGA) [76], the leading cancer-focused gene expression database, accessed via the UCSC Xena platform [28]. Additionally, we acquired essential domain knowledge from the Open Targets Platform [74] for gene-trait associations and the NCBI Genes database [10] for gene synonym mapping. Appendix E provides detailed information about our data acquisition process.
@@
Our team of 4 researchers designed the problem list and acquired relevant input data from public sources. During the pilot stage, a computational biologist and a doctoral student collaboratively developed example code for solving GTA problems related to two traits. By extracting common patterns from these examples, they developed an initial version of the analysis guidelines for the entire benchmark. Then they iteratively refined the guidelines based on their experience analyzing 193 more problems, establishing a working framework for benchmark curation.
@@
To evaluate annotation consistency, we measured Inter-Annotator Agreement (IAA) between the two annotation versions. Results demonstrate high quality, with an F1 score of 94.73% for dataset filtering. For more evaluation of the benchmark quality, we take the analysis results in the final benchmark as the reference set, and computed the average performance of the two annotation versions using the metrics defined in Section 3.3. This also serves as a strong baseline of human expert performance, with results presented in Table 2. The set of identified significant genes is sensitive to specific choices made during cohort-specific feature encoding, where multiple reasonable approaches often exist. Despite this inherent ambiguity in gene selection, the AUROC score of 0.89 demonstrates high consistency among annotators, validating the reliability of our benchmark.
@@ 3.3 Tasks and metrics
Gene expression data preprocessing involves complex decisions that affect both attribute names and sample IDs. We evaluate preprocessing performance using three complementary metrics: (i) Attribute Jaccard (AJ) is the Jaccard similarity between sets of attributes of two datasets. It evaluates how well the method extracts attributes from the dataset by encoding clinical features and normalizing gene symbols. (ii) Sample Jaccard (SJ) is the Jaccard similarity between sets of sample IDs of two datasets. It measures how well the method integrates features from the same samples and handles missing values. Based on these metrics, we define (iii) Composite Similarity Correlation (CSC) as the product of the Attribute Jaccard, Sample Jaccard, and the average Pearson correlation of common feature vectors (shared rows and columns) between datasets. This metric captures both structural and numerical similarity of datasets, so we consider it as the primary metric for evaluating preprocessing quality. We compute these scores per dataset and average across all datasets in our benchmark.
@@
Statistical analysis in our benchmark identifies trait-related significant genes, which we evaluate using multiple complementary metrics. We frame GTA analysis as a binary classification task, where each gene is classified as either significant or non-significant to a problem. We assess the classification performance using Precision, Recall, and F1 scores. To address the challenge of imbalanced gene sets, we employ the Area Under the Receiver Operating Characteristic curve (AUROC) [34], which provides a threshold-independent measure of discrimination ability. Additionally, we incorporate the Enrichment Score (ES) from Gene Set Enrichment Analysis (GSEA) [72], a standard bioinformatics metric that captures the biological relevance of identified gene sets. We designate AUROC as our primary metric due to its robustness to class imbalance and widespread acceptance in the machine learning community. All metrics are computed per problem and averaged across all problems in our benchmark.
@@ 4.1 Motivation and agent design
We believe that integrating these components is essential for enabling agent systems to effectively tackle the complex task of gene expression data analysis. Thus, to propose reasonable baselines for our benchmark, inspired by human bioinformaticians’ workflow in gene data analysis, we propose GenoAgent, a team of LLM-based agents with specialized roles mirroring those in a genomic data science team. A Project Manager coordinates the analysis process for solving each GTA problem, assigning tasks to agents with the standardized pipeline in our benchmark as instructions. Two programming agents, Data Engineer and Statistician, focus on the data preprocessing and statistical analysis tasks, respectively. When writing code to perform specific steps, they can access, adapt, and utilize function tools from a provided library file as needed. A Code Reviewer agent helps the programming agents debug code and verify that their code follows the instructions. A Domain Expert agent provides professional knowledge consultation to programming agents when required for data processing.
@@ 5 Experiment
This section describes our experiments to evaluate GenoAgent and other baseline methods on the GenoTEX benchmark. We conducted an end-to-end evaluation where methods process raw input data to complete the full analysis for solving GTA problems. Additionally, we assessed these methods on each individual task to identify specific strengths and weaknesses. All evaluations use the tasks and metrics defined in Section 3.3. We tested GenoAgent on a wide range of representative LLMs, as shown in Table 2.
@@
Table 2: End-to-end performance of GenoAgent on various LLMs; ablation results of GenoAgent with OpenAI o1 model; human expert performance, and simple data-free baselines. We report performance on the GTA problems in our benchmark, on the additional task of trait prediction, and runtime performance including success rate and efficiency metrics. Success rate (Succ.) refers to the percentage of problems for which the generated analysis code executes successfully. We use 00 as the trait prediction scores when code execution fails. “Tk.(K)” reports average input/output tokens (in thousands) per problem, while “Cost (\$)” and “Time (s)” represent average API cost and runtime per problem.
Model GTA Analysis Trait Prediction Runtime Performance
@@
Model GTA Analysis Trait Prediction Runtime Performance
@@ 5.1 Results
For comparison, we implemented a simple baseline where an OpenAI o1 model was directly prompted to identify significant genes for each problem without access to any datasets or additional information. This approach yielded poor performance (0.55 AUROC), underscoring the difficulty of this task. Beyond GTA analysis, we also measured the agents’ ability to predict binary traits using gene expression features and other covariates from their selected datasets, with strong results confirming the validity of the regression models they implemented.
@@ 5.2 Discussions
While our results demonstrate the potential of LLM-based agentic methods for gene expression data analysis, they also highlight important limitations of current approaches.